HUBBLE SPACE TELESCOPE DAILY REPORT #1970 PERIOD COVERED: 0000Z (UTC) 09/24/97 - 0000Z (UTC) 09/25/97 Daily Status Report as of 268/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed NIC/1/2/3 7421 (The Relation Between Quasar Luminosity and Host Galaxy Mass) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain ground-based, near-infrared imaging of nearby quasars which have indicated a relation between the luminosity of the nucleus and the luminous mass of the host galaxy. The observations completed as planned, with no reported problems. 1.4 Completed Two Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.5 Completed Four Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.6 Completed Three Sets of NIC/1/2/3 7793 (NICMOS MULTIACCUM Dark Characterization) The Near Infrared Camera and Multi-Object Spectrometer (NICMOS) was used to obtain on-orbit darks for the SCAMRR and STEP 256-sample sequences after the flight software change for the pedestal effect on day 233. The observations were completed as planned, and no anomalies were reported. 1.7 Completed STIS/CCD 7642 (CCD Red Light PSF Halo) The Space Telescope Imaging Spectrograph (CCD) was used to measure the PSF of a red star with the CCD at red wavelengths, to characterize the internal scattering of long-wavelength light within the CCD itself. The observations were completed as planned, with no reported problems. 1.8 Completed WF/PC-2/FOC/96 6820 (A Very Deep Luminosity Function for the Galactic Bulge) The WF/PC-2 and the FOC (f/96) were used to make very deep imaging observations of the nuclear bulge in galaxy NGC-6558-FL in order to study its luminosity function. This is the only old, metal-rich stellar population that we can study down to M-I~eq +12, unaffected by the dynamical processes that are known to occur in globular clusters. The observations were completed as scheduled, and no problems were reported. 1.9 Completed WF/PC-2 6431 (Determination of the Extragalactric Distance Scale) The WF/PC-2 was used to make observations of the galaxy NGC-1326A to be used in measuring the periods of Cepheid variable stars within this galaxy. This will allow for a more accurate estimate of the distances to these galaxies through the period-luminosity relation for the Cepheid stars. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of NIC/1/2/3 7729 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain "JHK" imaging of sky regions while another instrument was prime. These observations were part of the short NICMOS one-orbit parallel program. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Three Sets of STIS/CCD 7781 (STIS Non-Scripted Parallel Archive Proposal - Full Orbit) The Space Telescope Imaging Spectrograph (MA2) was used for a non-scripted parallel archive proposal (one full orbit). The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Three Sets of NIC/1/2 7222 (Search for the Main Sequence at the Galactic Center) The Near Infrared Camera (Cameras One and Two) was used to search for O and B stars for the various lines of evidence that indicate that the galactic center has formed stars within the last 10-200 million years. This will give a better age determination for the last epoch of star formation. The observations were completed without incident. 1.13 Completed Two Sets of NIC/1/2/3 7726 (NICMOS One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to make one orbit parallel calibration observations. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed NIC/1/2/3 7608 (NICMOS Focus Monitoring) The Near Infrared Camera (Cameras One, Two, and Three) was used to determine and monitor the optimal focus and tilt settings for all three NICMOS cameras. This proposal is built to run NIC1/2/3 sweeps on a 2-week basis. The observations completed without incident with no reported anomalies. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: The SSR interrupt log was dumped at 267/1250Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH